# Small Models, Big Insights: Leveraging Slim Proxy Models to Decide When and What to Retrieve for LLMs

### v2

@@ Abstract
The integration of large language models (LLMs) and search engines represents a significant evolution in knowledge acquisition methodologies. However, determining the knowledge that an LLM already possesses and the knowledge that requires the help of a search engine remains an unresolved issue. Most existing methods solve this problem through the results of preliminary answers or reasoning done by the LLM itself, but this incurs excessively high computational costs. This paper introduces a novel collaborative approach, namely SlimPLM, that detects missing knowledge in LLMs with a slim proxy model, to enhance the LLM’s knowledge acquisition process. We employ a proxy model which has far fewer parameters, and take its answers as heuristic answers. Heuristic answers are then utilized to predict the knowledge required to answer the user question, as well as the known and unknown knowledge within the LLM. We only conduct retrieval for the missing knowledge in questions that the LLM does not know. Extensive experimental results on five datasets with two LLMs demonstrate a notable improvement in the end-to-end performance of LLMs in question-answering tasks, achieving or surpassing current state-of-the-art models with lower LLM inference costs. [^1] Our code and datasets are available at https://github.com/plageon/SlimPlm
@@
[…]
**footnotetext: This work was done when Jiejun Tan was doing internship at Baichuan Intelligent Technology.footnotetext: Corresponding author.
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant prowess in various natural language processing (NLP) tasks Achiam et al. (2023), attributed to their advanced language comprehension and generation capabilities. Despite being trained on extensive text corpora, these models occasionally produce hallucinated content Zhou et al. (2020); Maynez et al. (2020). To tackle this problem, the integration of retrieval systems with LLMs has been proposed, enabling access to external knowledge bases for more accurate and reliable text generation.
Retrieval-augmented generation (RAG) involves using a retrieval system to supplement LLMs with relevant external information, thereby improving text generation quality (Peng et al., 2023; He et al., 2022). Yet, recent studies have suggested that retrieval may not always be beneficial. In cases where LLMs can adequately respond without external knowledge, retrieval may introduce irrelevant information, potentially degrading performance Kadavath et al. (2022); Wang et al. (2023b); Shi et al. (2023a); Petroni et al. (2020a). Therefore, it is critical to determine when retrieval is necessary for user questions Shuster et al. (2021). The challenge lies in identifying questions that exceed the LLMs’ intrinsic knowledge and require external retrieval, due to the prevalence of content hallucination. Efforts to address this challenge can be categorized into two groups: (1) The first group of methods involves fine-tuning LLMs for RAG scenarios, allowing them to autonomously signal the need for external knowledge Nakano et al. (2021); Liu et al. (2023b); Qin et al. (2023b). This method, while effective, demands substantial computational resources and risks diminishing the LLMs’ general capabilities due to potential catastrophic forgetting Kotha et al. (2023); Zhai et al. (2023). (2) The second category avoids direct tuning of LLMs, assessing the necessity for retrieval based on the quality of the generated content or specific indicators within it Ram et al. (2023); Min et al. (2022). However, this approach still has its drawbacks, as it requires multiple inferences, thereby increasing both the inference costs and the latency of responses to user questions.
@@ 2.1 Retrieval-augmented Generation (RAG)
Before the rise of LLMs, small language models Raffel et al. (2019); Lewis et al. (2019) possess limited knowledge, so retrieval-augmentation is used to assist knowledge-incentive QA tasks. Previous works often design a fixed module for retrieval-augmentation Sun et al. (2022); Lewis et al. (2020); Izacard and Grave (2020). Since the release of ChatGPT, LLMs have been developing rapidly Zhao et al. (2023). LLMs possess abundant intrinsic knowledge, but hallucination issue means they also have to rely on retrieval-augmentation to improve factuality Shuster et al. (2021); White (2023).
Since LLMs have a rather fixed architecture, retrieval-augmented LLMs generally leave LLMs’ architecture unchanged while pre-pending grounding documents to the input Ram et al. (2023); Yu et al. (2023b). Generally, a typical RAG pipeline Zhu et al. (2023); Liu et al. (2023a); Shi et al. (2023b) consists of a query rewriter Wang et al. (2023a); Gao et al. (2022), a retriever Guu et al. (2020); Neelakantan et al. (2022), a filter or reranker Yoran et al. (2023); Yu et al. (2023a); Xu et al. (2023), and an LLM as reader. Some methods attempt to enhance preposition components for overall performance improvement. Others fine-tune the LLM to fit into RAG tasks Asai et al. (2023); Kadavath et al. (2022).
@@ 2.2 Retrieval Necessity Judgment
Except for modifications to RAG pipeline components, retrieval necessity judgment is a more closely related topic. There are several streams to solve this issue.
(1) Fine-tuning LLMs is effective, but incurs significant computational costs Qin et al. (2023a); Lin et al. (2022). Some work fine-tune the LLM to imitate a human-like web-browsing behavior Schick et al. (2023); Nakano et al. (2021). Self-RAG Asai et al. (2023) introduces special tokens called reflection tokens to control retrieval behavior.
(2) Another intuitive approach involves assessing the LLM’s confidence based on logits produced by the model Jiang et al. (2020); Guo et al. (2017). FLARE Jiang et al. (2023) dynamically triggers RAG if logits fall below a certain threshold.
(3) Other work use iterative prompts to decide if extra information is required Wei et al. (2022); Liu et al. (2021); Rubin et al. (2021), or synergy Chain-of-Thought prompting Wei et al. (2022) with RAG Press et al. (2022); Khattab et al. (2022). For example, ReAct Yao et al. (2022) alternates between generating thoughts and actions, creating a series of thought-action-observation steps.
(4) Assessing the difficulty or popularity of user questions to judge retrieval necessity is also a viable approach Mallen et al. (2022). SKR Wang et al. (2023b) refers to similar questions they have previously encountered to judge retrieval necessity.
By contrast, SlimPLM conducts retrieval necessity judgment by assessing the answer generated by a smaller LLM. SlimPLM does not increase LLM inference times while improving judgment accuracy.
@@ 2.3 Query Formulation
Besides when to retrieve, what to retrieve is a closely related topic. Because it would be crude if search necessity of the user question gets judged. Retrieval necessity judgment work mentioned above also proposed query rewriting solutions. Many work prompt the LLM to generate queries itself Yao et al. (2022); Press et al. (2022); Schick et al. (2023). Some work take the formerly generated contents as queriesShao et al. (2023); Asai et al. (2023), or take a step further, mask the low logits tokens in the generated contents Jiang et al. (2023). Other work employ a specialy fine-tuned query rewriting model to rewrite user question or formerly generated contents Wang et al. (2023a); Ma et al. (2023).
By contrast, SlimPLM conducts query formulation by scrutinizing the answer generated by a smaller LLM providing an accurate perception of the necessary knowledge.
@@ 3.2 Retrieval Necessity Judgment
Input: <SYS> You are a helpful assistant. Your task is to parse user input into structured formats and accomplish the task according to the heuristic answer. </SYS> Heuristic answer: {Heuristic Answer} Question: {user question} Retrieval Necessity Judgment Output: Output: Known (True / False)
@@ 3.3 Retrieval Target Determination
To train the query rewriting model, we collect and annotate a dataset with the help of GPT-4Achiam et al. (2023). In each dataset used in our experiments, we sample 1,000 user questions. We utilize the method of instruction fine-tuning Ouyang et al. (2022); Chung et al. (2022) to fine-tune a decoder-only generative model, accomplishing the task of claim extraction and query rewriting in a single round. Our instructions and the model output are displayed as follows.
Input: <SYS> You are a helpful assistant. Your task is to parse user input into structured formats and accomplish the task according to the heuristic answer. </SYS> Heuristic answer: {Heuristic Answer} Question: {User Question} Query Rewrite Output: Output: <Claim> Claim 1 <Query> Query 1 <Claim> Claim 2 <Query> Query 2, ...
@@ Appendix C Case Study
reference: Åre and Östersund, Sweden will host the next World Winter Games between February 2 to 13, 2021. It will mark the first time that Sweden has ever hosted the Special Olympics World Games.
Error Type: irrelevant reference
reference: The EOC launched the bid process on 20 September 2018 after a meeting of the constituent National Olympic Committees in Stockholm.
Error Type: irrelevant reference
@@
reference: The 2017 NBA playoffs began on April 15, 2017. It concluded with the Golden State Warriors defeating the Cleveland Cavaliers 4 games to 1 in the NBA Finals, their third consecutive meeting at the Finals.
Error Type: incomplete reference
reference: This Finals was the first time in NBA history the same two teams had met for a third consecutive year. The Cavaliers sought to repeat as champions after winning the championship in 2016, while the Warriors won the first meeting in 2015.
Error Type: incomplete reference
@@
Your task is to perform text analysis on user conversations, and complete the last json item. You need to follow the following rules:
1. Classify user conversations into the following categories: text rewriting, mathematical problems, knowledge questions, text creation, table processing, translation, summarization, logical reasoning, open qa, coding, text classification, information extraction, brainstorming, exams, role-playing, others. The format should be a string and stored in the task field.
2. Determine whether the answer of user input is closely related to current datetime, and store it in the timeliness field in boolean format.
3. If the user’s request involves reasoning, each reasoning process should be described as questions and split into as many sub-questions as possible.
4. The sub-questions after splitting should be placed in the question field in questions, and the sub-questions should be fully described without using pronouns such as “he”, “this”, or “that”.
5. If the sub-question involves very strict factual information such as personal relationships, time, location, policies, regulations, etc., which requires the use of a search engine to answer, then it needs to be marked as needSearch=true, and the generated search term should be placed in searchWord.
6. If the sub-question is a chit-chat question such as "how are you" or a pure mathematical problem, coding, logical reasoning, creative thinking, or common sense problem, then no search is needed.
7. Extract the entities and events involved in the user’s request and store them in the entities and events fields respectively. The format is a list of strings. Note that the entities and events should be higly informative, and should not be a user instruction or a question.
@@
<<SYS>>
Now, based on the following reference and your knowledge, please answer the question more succinctly and professionally. The reference is delimited by triple brackets [[[]]]. The question is delimited by triple parentheses ((())). You should include as many possible answers as you can.
<</SYS>>
Reference: [[[reference]]],
question: (((question)))
@@
<<SYS>>
Now, based on the following reference and your knowledge, please answer the question more succinctly and professionally. The reference is delimited by triple brackets [[[]]]. The question is delimited by triple parentheses ((())). You are not allowed to add fabrications or hallucinations.
<</SYS>>
Reference: [[[reference]]],
question: (((question)))
@@
<<SYS>>
Given the following question, answer it by providing follow up questions and intermediate answers. If no follow up questions are necessary, answer the question directly.
<<SYS>>
Question: Who lived longer, Muhammad Ali or Alan Turing?
Are follow up questions needed here: Yes.
Follow up: How old was Muhammad Ali when he died?
Intermediate answer: Muhammad Ali was 74 years old when he died.
Follow up: How old was Alan Turing when he died?
Intermediate answer: Alan Turing was 41 years old when he died.
So the final answer is: Muhammad Ali
Question: When was the founder of craigslist born?
Are follow up questions needed here: Yes.
Follow up: Who was the founder of craigslist?
Intermediate answer: Craigslist was founded by Craig Newmark.
Follow up: When was Craig Newmark born?
Intermediate answer: Craig Newmark was born on December 6, 1952.
So the final answer is: December 6, 1952
Question: question
@@
Given the following reference, answer it by a brief sentence. You are not allowed to add fabrications or hallucinations.
reference
Question: How old was Muhammad Ali when he died?
Answer: Muhammad Ali was 74 years old when he died.
Question: Who was the founder of craigslist?
Answer: Craigslist was founded by Craig Newmark.
Question: Who was the father of Mary Ball Washington?
Answer: The father of Mary Ball Washington was Joseph Ball.
Question: Who is the director of Casino Royale?
Answer: The director of Casino Royale is Martin Campbell.
Question: question
Answer:

### v3

@@ Abstract
The integration of large language models (LLMs) and search engines represents a significant evolution in knowledge acquisition methodologies. However, determining the knowledge that an LLM already possesses and the knowledge that requires the help of a search engine remains an unresolved issue. Most existing methods solve this problem through the results of preliminary answers or reasoning done by the LLM itself, but this incurs excessively high computational costs. This paper introduces a novel collaborative approach, namely SlimPLM, that detects missing knowledge in LLMs with a slim proxy model, to enhance the LLM’s knowledge acquisition process. We employ a proxy model which has far fewer parameters, and take its answers as heuristic answers. Heuristic answers are then utilized to predict the knowledge required to answer the user question, as well as the known and unknown knowledge within the LLM. We only conduct retrieval for the missing knowledge in questions that the LLM does not know. Extensive experimental results on five datasets with two LLMs demonstrate a notable improvement in the end-to-end performance of LLMs in question-answering tasks, achieving or surpassing current state-of-the-art models with lower LLM inference costs. [^1] Our code and datasets are available at https://github.com/plageon/SlimPlm.
@@
Jiejun Tan1thanks: This work was done when Jiejun Tan was doing internship at Baichuan Intelligent Technology., Zhicheng Dou1thanks: Corresponding author., Yutao Zhu1, Peidong Guo2 Kun Fang2,  and Ji-Rong Wen1 1Gaoling School of Artificial Intelligence, Renmin University of China 2 Baichuan Intelligent Technology {zstanjj, dou}@ruc.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated significant prowess in various natural language processing (NLP) tasks OpenAI (2023), attributed to their advanced language comprehension and generation capabilities. Despite being trained on extensive text corpora, these models occasionally produce hallucinated content Zhou et al. (2021); Maynez et al. (2020). To tackle this problem, the integration of retrieval systems with LLMs has been proposed, enabling access to external knowledge bases for more accurate and reliable text generation.
Retrieval-augmented generation (RAG) involves using a retrieval system to supplement LLMs with relevant external information, thereby improving text generation quality (Peng et al., 2023; He et al., 2023). Yet, recent studies have suggested that retrieval may not always be beneficial. In cases where LLMs can adequately respond without external knowledge, retrieval may introduce irrelevant information, potentially degrading performance Kadavath et al. (2022); Wang et al. (2023b); Shi et al. (2023a); Petroni et al. (2020). Therefore, it is critical to determine when retrieval is necessary for user questions Shuster et al. (2021). The challenge lies in identifying questions that exceed the LLMs’ intrinsic knowledge and require external retrieval, due to the prevalence of content hallucination. Efforts to address this challenge can be categorized into two groups: (1) The first group of methods involves fine-tuning LLMs for RAG scenarios, allowing them to autonomously signal the need for external knowledge Nakano et al. (2021); Liu et al. (2023b); Qin et al. (2023b). This method, while effective, demands substantial computational resources and risks diminishing the LLMs’ general capabilities due to potential catastrophic forgetting Kotha et al. (2023); Zhai et al. (2023). (2) The second category avoids direct tuning of LLMs, assessing the necessity for retrieval based on the quality of the generated content or specific indicators within it Ram et al. (2023); Min et al. (2022). However, this approach still has its drawbacks, as it requires multiple inferences, thereby increasing both the inference costs and the latency of responses to user questions.
@@ 2.1 Retrieval-Augmented Generation (RAG)
RAG has been studied for a long time. In the era of pre-trained language models, RAG has been applied to provide models with relevant knowledge, significantly enhancing the generation quality in applications such as dialogue systems Tahami et al. (2020); Tao et al. (2019) and question-answering systems Izacard and Grave (2021); Tahami et al. (2020). With the development of LLMs, RAG has emerged as a crucial strategy to tackle the problem of hallucination and outdated information Shuster et al. (2021); White (2023).
The mainstream RAG methods follow a “retrieve-then-read” architecture. In this setup, a retrieval module first gathers external knowledge, providing additional context that is subsequently processed by LLMs to generate the final output Ram et al. (2023); Yu et al. (2023b). Typically, a RAG pipeline Zhu et al. (2023); Liu et al. (2023a); Shi et al. (2023b) includes several components: a query rewriter that refines the initial query Wang et al. (2023a); Gao et al. (2023), a retriever that fetches relevant documents Guu et al. (2020); Neelakantan et al. (2022), a filter or reranker Yoran et al. (2023); Yu et al. (2023a); Xu et al. (2023) that ensures only the most relevant knowledge is kept, and an LLM as reader that generates the final results. To optimize these systems, some approaches focus on enhancing individual components of the RAG architecture to improve overall performance Zhu et al. (2024); Jin et al. (2024), while others involve direct fine-tuning of the LLM to better integrate with RAG-specific tasks Asai et al. (2023); Kadavath et al. (2022).
@@ 2.2 Retrieval Necessity Judgment
In a Retrieval-Augmented Generation (RAG) system, a critical challenge is determining when to initiate the retrieval process. Several approaches have been proposed to address this issue:
(1) Fine-tuning Large Language Models (LLMs) has proven effective but comes with substantial computational costs Qin et al. (2023a); Lin et al. (2022). Some studies have focused on fine-tuning the LLM to mimic human-like web browsing behavior Schick et al. (2023); Nakano et al. (2021). Self-RAG Asai et al. (2023) introduces special tokens known as reflection tokens to regulate retrieval behavior.
(2) Another intuitive approach involves evaluating the LLM’s confidence based on the logits generated by the model Jiang et al. (2021); Guo et al. (2017). FLARE Jiang et al. (2023) dynamically activates RAG if the logits fall below a predefined threshold.
(3) Other research has employed iterative prompting to determine if additional information is required Wei et al. (2022); Liu et al. (2022); Rubin et al. (2022), or has combined Chain-of-Thought prompting Wei et al. (2022) with RAG Press et al. (2023); Khattab et al. (2022). For instance, ReAct Yao et al. (2023) alternates between generating thoughts and actions, creating a sequence of thought-action-observation steps.
(4) Evaluating the complexity or popularity of user questions to assess the need for retrieval is also a feasible approach Mallen et al. (2023). SKR Wang et al. (2023b) refers to similar questions it has previously encountered to determine the necessity of retrieval.
Distinct from these existing methods, SlimPLM evaluates the necessity of retrieval by analyzing the answer generated by a smaller LLM. This approach does not increase LLM inference times while enhancing judgment accuracy.
@@ 2.3 Query Formulation
In addition to determining when to retrieve information, the question of what to retrieve is also of great importance. A simplistic approach that merely judges the necessity of retrieval based on the user’s query would be inadequate. The aforementioned retrieval necessity judgment work has also proposed solutions for query rewriting. Numerous studies have encouraged the use of Large Language Models (LLMs) to autonomously generate queries Yao et al. (2023); Press et al. (2023); Schick et al. (2023). Some research has utilized previously generated content as queries Shao et al. (2023); Asai et al. (2023), or have taken a further step by masking low logit tokens within the generated content Jiang et al. (2023). Other studies have employed specially fine-tuned query rewriting models to rewrite either the user’s question or previously generated content Wang et al. (2023a); Ma et al. (2023).
In contrast, SlimPLM formulates queries by meticulously analyzing the answers generated by a smaller LLM, thereby providing an accurate understanding of the required knowledge.
@@ 3.2 Retrieval Necessity Judgment
{mdframed}
[backgroundcolor=gray!5, roundcorner=3pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt] Input:
<SYS> You are a helpful assistant. Your task is to parse user input into structured formats and accomplish the task according to the heuristic answer. </SYS>
Heuristic answer: {Heuristic Answer}
Question: {user question}
Retrieval Necessity Judgment Output:
Output:
Known (True / False)
@@ 3.3 Retrieval Target Determination
To train the query rewriting model, we collect and annotate a dataset with the help of GPT-4OpenAI (2023). In each dataset used in our experiments, we sample 1,000 user questions. We utilize the method of instruction fine-tuning Ouyang et al. (2022); Chung et al. (2022) to fine-tune a decoder-only generative model, accomplishing the task of claim extraction and query rewriting in a single round. Our instructions and the model output are displayed as follows.
{mdframed}
[backgroundcolor=gray!5, roundcorner=3pt, innerleftmargin=10pt, innerrightmargin=10pt, innertopmargin=10pt, innerbottommargin=10pt] Input:
<SYS> You are a helpful assistant. Your task is to parse user input into structured formats and accomplish the task according to the heuristic answer. </SYS>
Heuristic answer: {Heuristic Answer}
Question: {User Question}
Query Rewrite Output:
Output:
<Claim> Claim 1 <Query> Query 1 <Claim> Claim 2 <Query> Query 2, ...
@@ 5 Conclusion
Acknowledgments
This work was supported by Beijing Natural Science Foundation No. L233008, CCF-BaiChuan-Ebtech Foundation Model Fund, National Natural Science Foundation of China No. 62272467, the fund for building world-class universities (disciplines) of Renmin University of China, and Public Computing Cloud, Renmin University of China. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.
@@ Appendix C Case Study
reference: Åre and Östersund, Sweden will host the next World Winter Games between February 2 to 13, 2021. It will mark the first time that Sweden has ever hosted the Special Olympics World Games. Error Type: irrelevant reference reference: The EOC launched the bid process on 20 September 2018 after a meeting of the constituent National Olympic Committees in Stockholm. Error Type: irrelevant reference
@@
reference: The 2017 NBA playoffs began on April 15, 2017. It concluded with the Golden State Warriors defeating the Cleveland Cavaliers 4 games to 1 in the NBA Finals, their third consecutive meeting at the Finals. Error Type: incomplete reference reference: This Finals was the first time in NBA history the same two teams had met for a third consecutive year. The Cavaliers sought to repeat as champions after winning the championship in 2016, while the Warriors won the first meeting in 2015. Error Type: incomplete reference
@@
Your task is to perform text analysis on user conversations, and complete the last json item. You need to follow the following rules: 1. Classify user conversations into the following categories: text rewriting, mathematical problems, knowledge questions, text creation, table processing, translation, summarization, logical reasoning, open qa, coding, text classification, information extraction, brainstorming, exams, role-playing, others. The format should be a string and stored in the task field. 2. Determine whether the answer of user input is closely related to current datetime, and store it in the timeliness field in boolean format. 3. If the user’s request involves reasoning, each reasoning process should be described as questions and split into as many sub-questions as possible. 4. The sub-questions after splitting should be placed in the question field in questions, and the sub-questions should be fully described without using pronouns such as “he”, “this”, or “that”. 5. If the sub-question involves very strict factual information such as personal relationships, time, location, policies, regulations, etc., which requires the use of a search engine to answer, then it needs to be marked as needSearch=true, and the generated search term should be placed in searchWord. 6. If the sub-question is a chit-chat question such as "how are you" or a pure mathematical problem, coding, logical reasoning, creative thinking, or common sense problem, then no search is needed. 7. Extract the entities and events involved in the user’s request and store them in the entities and events fields respectively. The format is a list of strings. Note that the entities and events should be higly informative, and should not be a user instruction or a question.
@@
<<SYS>> Now, based on the following reference and your knowledge, please answer the question more succinctly and professionally. The reference is delimited by triple brackets [[[]]]. The question is delimited by triple parentheses ((())). You should include as many possible answers as you can. <</SYS>> Reference: [[[reference]]], question: (((question)))
@@
<<SYS>> Now, based on the following reference and your knowledge, please answer the question more succinctly and professionally. The reference is delimited by triple brackets [[[]]]. The question is delimited by triple parentheses ((())). You are not allowed to add fabrications or hallucinations. <</SYS>> Reference: [[[reference]]], question: (((question)))
@@
<<SYS>> Given the following question, answer it by providing follow up questions and intermediate answers. If no follow up questions are necessary, answer the question directly. <<SYS>> Question: Who lived longer, Muhammad Ali or Alan Turing? Are follow up questions needed here: Yes. Follow up: How old was Muhammad Ali when he died? Intermediate answer: Muhammad Ali was 74 years old when he died. Follow up: How old was Alan Turing when he died? Intermediate answer: Alan Turing was 41 years old when he died. So the final answer is: Muhammad Ali Question: When was the founder of craigslist born? Are follow up questions needed here: Yes. Follow up: Who was the founder of craigslist? Intermediate answer: Craigslist was founded by Craig Newmark. Follow up: When was Craig Newmark born? Intermediate answer: Craig Newmark was born on December 6, 1952. So the final answer is: December 6, 1952 Question: question
@@
Given the following reference, answer it by a brief sentence. You are not allowed to add fabrications or hallucinations. reference Question: How old was Muhammad Ali when he died? Answer: Muhammad Ali was 74 years old when he died. Question: Who was the founder of craigslist? Answer: Craigslist was founded by Craig Newmark. Question: Who was the father of Mary Ball Washington? Answer: The father of Mary Ball Washington was Joseph Ball. Question: Who is the director of Casino Royale? Answer: The director of Casino Royale is Martin Campbell. Question: question Answer: